Hello,
3. Find the absolute inequality corresponding to x <= -22/3 or x >= 11/3
I came up with an answer and was tould to work on it alittle more. Could someone please show me the correct steps to find the correct answer. Thank you!

Answer by stanbon(57354)

Find the absolute inequality corresponding to x <= -22/3 or x >= 11/3

Draw a number line and put -22/3 and 11/3 on it.
Find the point half-way between the two points as follows:
[11/3 + (-22/3)]/2 = -11/6
Find the distance from this mid-point to the end point 11/3:
11/3-(-11/6)=22/6+11/6= 33/6
With an absolute value statement you need to describe all "x-values"
that are > or = to 33/6 away from -11/6.
That is |x+(11/6)|>=33/6
Cheers,
stan H.